OpenStudy (anonymous):

Evaluate to four decimal places Please explain how you found your answer cot183

OpenStudy (campbell_st):

find tan(183) which is 3rd quadrant, positive and has the same value as tan(3) (3rd quadrant angles are 180 + x) when you know tan (3) take the reciprocal to find cot (183) which is the same as cot(3) hope this makes sense.

OpenStudy (anonymous):

what reciprocal?

OpenStudy (anonymous):

and what is tan 3? and what does this all mean??? i'm so confused

OpenStudy (campbell_st):

turn it up side down so if you have 2/3 the reciprocal is 3/2 you calculator may have a key \[x^{-1} \] when you get the answer for tan(3) press the button of find \[\frac{1}{\tan(3)}\]

OpenStudy (anonymous):

but why tan3? Where do you get that?

OpenStudy (campbell_st):

its a 3rd quadrant angle and tan is positive, so cot is positive. the angle is the same as (180 + x) in this question x = 3 so you are really finding cot(3) = 1/tan(3)

OpenStudy (anonymous):

is there another method of doing that? like how do you put it in your calculator?

OpenStudy (campbell_st):

ok... so have you learnt about angles in any quadrant 

OpenStudy (campbell_st):

yep the easy way to do it is to graph cot(x) then look at the value where x = 183

OpenStudy (anonymous):

...how do you graph that on a ti-84 ?? please help

OpenStudy (campbell_st):

ok... so do you have a y = key under the screen on the left...?

OpenStudy (anonymous):

yes

OpenStudy (campbell_st):

ok... press it

OpenStudy (anonymous):

alright

OpenStudy (campbell_st):

so the cursor is next to the 1st equals sign..? type 1/tan(x) press the tan key and in the brackets just press x I think you have an (x, r Theta) button.

OpenStudy (anonymous):

got it

OpenStudy (campbell_st):

once you have entered it press enter... you maky have to press the window key to change the graph..

OpenStudy (campbell_st):

set the x values to min180, max 270 scale 50

OpenStudy (campbell_st):

when you have the graph... press the trace key, at the top under the screen and enter 183... it will show a value.

OpenStudy (campbell_st):

good luck

OpenStudy (anonymous):

sweet I got it! THank you very much :)!

OpenStudy (anonymous):

do you always set your x values to min180, max 270 scale 50?

OpenStudy (campbell_st):

the angle you were looking at was 183 which is between 180 and 270 set the x value to either side of your desired value... graphing in units of 90 degrees give you the opportunity to see a graph...

OpenStudy (campbell_st):

you can adjust the scale to something smaller... if you wish